OpenStudy (anonymous):

calculus question. Totally have no idea what Im doing... well actually I do but I suck. calculus question. Totally have no idea what Im doing... well actually I do but I suck. @Mathematics

OpenStudy (anonymous):

gimmie a second to type this up

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi \over 2}\int\limits_{0}^{3}\int\limits_{0}^{e ^{-r^2}}r \space dz \space dr \space d \theta\]sketch the the solid region whose volume is given by the iterated integral and evaluate the iterated integral

OpenStudy (anonymous):

I know I let u = -r^2 and that is about it lol

OpenStudy (anonymous):

i think I can also let du=-2rdr

OpenStudy (anonymous):

okay after this I have no clue. And Im not even sure if my substitution for u is right in the first place >.>

OpenStudy (eyust707):

always sketch it first

OpenStudy (eyust707):

the integral all the way to the right corelates to the first d?? we see at the end

OpenStudy (eyust707):

that means the first integral evaulates from 0 to e^-rz with respect to z

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

if you evaulaute the first z integral you get er-r^2

OpenStudy (eyust707):

i believe this is the function that represent the height of your cylinder

OpenStudy (eyust707):

now we can look at the bounds of r and thet to find the overall shape

OpenStudy (jamesj):

Here's the region for you:

OpenStudy (jamesj):

although that should be not e^(-9/2), but e^(-9)

OpenStudy (eyust707):